OpenStudy (anonymous):

Compare the fractions 4/7 1/6

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (anonymous):

Another way is by comparing decimals. \[\frac{4}{7} = 0.57143\] \[\frac{1}{6} = 0.16667\] 0.57143 > 0.16667 So by that: \[\frac{4}{7}>\frac{1}{6}\]